# LLM-A*: Large Language Model Enhanced Incremental Heuristic Search on Path Planning

### v1

@@ 1 Introduction
Existing path planning algorithms are capable of effectively completing planning tasks and ensuring the validity of their paths. However, as the environment and map scale up, algorithms like A* and its variants Hart et al. (1968b); Korf et al. (2001); Harabor and Grastien (2011); Jansen and Buro (2007) encounter an exponential increase in computational and memory demands. This occurs because the pathfinding process can become sub-optimal (see Figure 1), where the algorithm might spend unnecessary effort exploring less relevant areas, leading to exponential increases in time complexity as the map size enlarges.
@@
In this work, we propose LLM-A*, a new LLM based route planning method that synergizes the traditional A* algorithm with the global insights from Large Language Models. As illustrated in Fig. 1, this hybrid approach leverages LLM-generated waypoints to guide the path searching process, significantly reducing computational and memory costs. In addition, by integrating the standard L2 distance-based heuristic of A* with new heuristic values derived from these waypoints, LLM-A* addresses the granularity issues in LLM-generated solutions, ensuring the validity of the output paths.
We conducted extensive experiments across various environment to compare the performance of A* and LLM-A* (integrating LLAMA3 with few-shot prompting). As illustrated in Figure 3, A* exhibits exponential growth in both computational operations and storage requirements with linearly increasing environment scale. In contrast, LLM-A* shows a nearly linear growth pattern, indicating superior scalability. This suggests that LLM-A* is significantly more efficient in terms of both computation and memory, making it better suited for larger environments. Furthermore, our primary experimental results, summarized in Table 1, reveal that LLM-A* not only excels in scalability but also outperforms A* in baseline computational and memory efficiency. LLM-A* achieves significantly lower operation and storage ratios compared to A*, requiring less than about half the operations and storage needed by A* on average for the pathfinding process, thereby offering a robust and efficient solution for large-scale path planning.
@@ 2 Related Work
1:Input: START state $s_{0}$, GOAL state $s_{g}$, OBSTACLE state $obs$, heuristic function $h$, cost function $g$, Large Language Model $llm$
2:Output: Path $P$ from $s_{0}$ to $s_{g}$
3:Initialize the OPEN list $O=\{s_{0}\}$, CLOSE list $C=\{\}$, TARGET list $T=llm(s_{0},s_{g},obs)$, TARGET state $t=T.start$, $g(s_{0})=0$, $f(s_{0})=h(s_{0})$, $P=\{\}$
4:while $O\neq\emptyset$ do
5:     $s_{a}\leftarrow$ state in $O$ with the lowest $f$-cost
6:     if $s_{a}=s_{g}$ then
7:         return reconstruct_path($s_{a}$)
8:     Remove $s_{a}$ from $O$
9:     Add $s_{a}$ to $C$
10:     for all neighbors $s_{n}$ of $s_{a}$ do
11:         if $s_{n}\in C$ then
12:              continue
13:         if $s_{n}=t$ and $s_{g}\neq t$ then
14:              $t=T.next$
15:              update $f$-cost of $s$ in $O$
16:         Tentative cost $g_{tent}=g(s_{a})+cost(s_{a},s_{n})$
17:         if $s_{n}\notin O$ or $g_{tent}<g(s_{n})$ then
18:              Update path to $s_{n}$ to go through $s_{a}$
19:              $g(s_{n})=g_{tent}$
20:              $f(s_{n})=g(s_{n})+h(s_{n})+cost(t,s_{n})$
21:              if $s_{n}\notin O$ then
22:                  Add $s_{n}$ to $O$
23:return failure
Figure 2: LLM-A* Algorithm Pseudocode
@@ 3.2 LLM-A* Algorithm
LLM-A* integrates the global insights provided by Large Language Models (LLMs) with the A* algorithm’s optimal local search mechanism, where achieves a balance between the efficiency of the pathfinding process and optimality. The pseudocode for LLM-A* is shown in Figure 2, and it closely resembles the original A* algorithm.
@@
The pathfinding process of LLM-A* is similar to that of A*. It uses a heuristic function $h$, a cost function $g$, an OPEN list $O$, and a CLOSED list $C$. The algorithm searches through each state in $O$ until the goal state $s_{g}$ is reached. Each explored state $s_{a}$ is saved into $C$ to avoid redundant searches. The distinction that encapsulates the main differences between LLM-A* and A* happens during the expansion of the neighbor state $s_{n}$ (see in Figure 2:$13$-$15$). For each $s_{n}$, we check if it matches the current target $t$ from $T$. If the current $t$ is reached, $t$ is updated to the next target in $T$. Subsequently, the $f$-cost of every state in the current $O$ is re-computed, where the $f$-cost in LLM-A* is computed as the sum of the state’s cost, the heuristic value, and the cost from the state to current $t$ (see in Figure 2:$20$), defined as $f(s)=g(s)+h(s)+cost(t,s)$. This step introduces an additional computational amount to the pathfinding process, and the time complexity scales linearly with both the length of $T$ and the increasing size of $O$. However, it is important that this re-computation process ensures that the $f$-cost of visited states in $O$ remains accurate and updated with the new target $t$.
@@ 4.1 Dataset
These parameters define the structure and constraints of each map, ensuring consistency and relevance to the standard experimental environment conditions for search-based algorithms. Meanwhile, the map environment is able to scale properly for scalability experiment.
Methodology Base Model Prompt Approach Operation Ratio $\downarrow$ (%) Storage Ratio $\downarrow$ (%) Relative Path Length $\downarrow$ (%) Valid Path Ratio $\uparrow$ (%) A* - - 100 100 100 100 LLM GPT-3.5 Few-Shot - - 119.38 12.80 CoT - - 151.73 15.20 RePE - - 183.87 7.80 LLAMA3 Few-Shot - - 111.05 12.60 CoT - - 114.89 12.00 RePE - - 138.32 16.40 LLM-A* (Ours) GPT-3.5 Few-Shot 57.39 74.96 102.44 100 CoT 69.50 83.65 102.54 100 RePE 85.47 96.53 102.41 100 LLAMA3 Few-Shot 44.59 64.02 102.47 100 CoT 47.60 66.27 102.46 100 RePE 64.08 80.19 102.54 100
Table 1: Quantitative analysis of three pathfinding methodologies: the classical A* algorithm, an LLM-only approach, and our proposed LLM-A* approach. The methodologies are evaluated on the map size ($50\times 30$) of original samples. The LLM-only approaches explore the path without explicitly searching the space grid by grid, so we do not report the operation and storage ratio. The table includes the results from GPT-3.5 and LLAMA3 models with three prompting approaches: Few-Shot, Chain of Thought (CoT), and Recursive Path Evaluation (RePE) for both LLM-only and LLM-A* approaches (see Section 4.4 for details).
@@ 4.2 Experimental Setup
Our system allows search-based pathfinding in a continuous environment with modules for environment management, agent control, and visualization (see Section 4.1).
  • Environment Management: Configures the environment and provides feedback.
  • Agent Control: Customizes the agent’s operations using the algorithm and model.
  • Visualization: Offers real-time and final visual outputs for analysis.
@@ 4.4 Quantitative Analysis
Notably, the Recursive Path Evaluation (RePE) prompting method achieves the smallest increases in relative path length in LLM-A*, with increments of $2.41\%$ for the GPT-3.5 models, respectively. This suggests that RePE’s step-by-step progression and intrinsic reasoning capabilities improve the models’ ability to generate more optimal waypoints, resulting in more efficient paths. However, RePE underperforms compared to Chain of Thought (CoT) and few-shot prompting when used in the LLM-only approach. This indicates limitations in LLMs’ ability to execute end-to-end path planning and spatial-temporal reasoning, which not only affects their proficiency in sequentially reasoning out detailed path sequences but also leads to issues such as hallucinations and misunderstandings. These limitations can cause the model to generate incorrect or implausible paths, undermining the effectiveness of LLMs in isolated path planning tasks.
@@ 4.5 Qualitative Analysis
From the visualization in Figure 1, LLM-A* identifies the optimal path with only 140 operations, less than one-fifth the 859 operations required by A*, as well as the storage reduction. Both algorithms utilize a priority queue that stores the $f$-cost of each reached state, with the state having the lowest $f$-cost selected for exploration. The fundamental distinction between the two algorithms lies in their calculation of the $f$-cost or heuristic values.
@@ 5 Conclusion
Although around 90% of the paths generated by LLM-A* are optimal, our algorithm does not guarantee optimal path. While these cases are relatively few, they indicate that the algorithm may sometimes yield paths that are not the shortest or most efficient. Future improvements could focus on enhancing the optimality of the generated paths to ensure more consistent performance. Our experiments mainly utilized GPT-3.5-TURBO and LLAMA3-8B-16bit with basic prompt techniques. Although these models and prompts were adequate to validate the robustness of the LLM-A* algorithm, we did not explore a wider array of models or advanced prompt engineering strategies. Further testing with additional models and varied prompting methods could provide more comprehensive insights into the algorithm’s performance across different scenarios.

### v2

@@ 1 Introduction
Existing path planning algorithms are capable of effectively completing planning tasks and ensuring the validity of their paths. However, as the environment and map scale up, algorithms like A* and its variants Hart et al. (1968b); Korf et al. (2001); Harabor and Grastien (2011); Jansen and Buro (2007) encounter an exponential increase in computational and memory demands. This occurs because the pathfinding process can become sub-optimal (see Figure 1 and 2), where the algorithm might spend unnecessary effort exploring less relevant areas, leading to exponential increases in time complexity as the map size enlarges.
@@
Figure 2: Visual comparison of pathfinding efficiency Between A* and LLM-A*. This figure illustrates the performance differences between the traditional A* algorithm (left and upper images) and the LLM-A* algorithm (right and lower images). Red lines indicate the computed paths, blue dots mark the starting state, green dots indicate the goal state, gray areas represent visited states, and black lines denote obstacles. The LLM-A* algorithm demonstrates more efficient pathfinding by requiring significantly fewer visited states than A*.
@@ 2 Related Work
Recent work, such as the LLM3 framework Wang et al. (2024) leverages pre-trained LLMs to integrate symbolic task planning with continuous motion generation through motion failure reasoning, where LLM3 iteratively refines both symbolic actions and continuous parameters, significantly improving planning efficiency in dynamic environments, which aligns with our focus on LLMs’ adaptability in correcting low-level planning errors and enhancing resilience in dynamic conditions.
@@
1:Require: START state $s_{0}$, GOAL state $s_{g}$, OBSTACLE set $obs$, heuristic function $h()$, cost function $g()$, Large Language Model $llm()$
2:OPEN list $\mathcal{O}_{\text{open}}\leftarrow\{s_{0}\}$, CLOSE list $\mathcal{C}_{\text{close}}\leftarrow\{\}$, TARGET list $\mathcal{T}\leftarrow llm(s_{0},s_{g},\mathcal{O})$, TARGET state $t\leftarrow\mathcal{T}_{\text{start}}$, $g(s_{0})\leftarrow 0$, $f(s_{0})\leftarrow h(s_{0})$, $P\leftarrow\{\}$
3:while $\mathcal{O}_{\text{open}}\neq\emptyset$ do
4:     $s_{a}\leftarrow\arg\min_{s\in\mathcal{O}_{\text{open}}}f(s)$
5:     if $s_{a}=s_{g}$ then
6:         return reconstruct_path($s_{a}$)
7:     Remove $s_{a}$ from $\mathcal{O}_{\text{open}}$
8:     Add $s_{a}$ to $\mathcal{C}_{\text{close}}$
9:     for all neighbors $s_{n}$ of $s_{a}$ do
10:         if $s_{n}=t$ and $s_{g}\neq t$ then
11:              $t\leftarrow\mathcal{T}_{\text{next}}$
12:              Update $f$-cost of states in $\mathcal{O}_{\text{open}}$
13:         if $s_{n}\in\left(\mathcal{C}_{\text{close}}\cup\text{obs}\right)$ then
14:              continue
15:         Tentative cost $g_{\text{tent}}\leftarrow g(s_{a})+\text{cost}(s_{a},s_{n})$
16:         if $s_{n}\notin\mathcal{O}_{\text{open}}$ or $g_{\text{tent}}<g(s_{n})$ then
17:              Update path to $s_{n}$ to go through $s_{a}$
18:              $g(s_{n})\leftarrow g_{\text{tent}}$
19:              $f(s_{n})\leftarrow g(s_{n})+h(s_{n})+\text{cost}(t,s_{n})$
20:              if $s_{n}\notin\mathcal{O}_{\text{open}}$ then
21:                  Add $s_{n}$ to $\mathcal{O}_{\text{open}}$
22:return failure
@@ 3.2 LLM-A* Algorithm
LLM-A* integrates the global insights provided by Large Language Models (LLMs) with the A* algorithm’s optimal local search mechanism, where achieves a balance between the efficiency of the pathfinding process and optimality. The pseudocode for LLM-A* is shown in Algorithm 1, and it closely resembles the original A* algorithm.
@@
The pathfinding process of LLM-A* is similar to that of A*. It uses a heuristic function $h$, a cost function $g$, an OPEN list $O$, and a CLOSED list $C$. The algorithm searches through each state in $O$ until the goal state $s_{g}$ is reached. Each explored state $s_{a}$ is saved into $C$ to avoid redundant searches. The distinction that encapsulates the main differences between LLM-A* and A* happens during the expansion of the neighbor state $s_{n}$ (see in Algorithm 1:$13$-$15$). For each $s_{n}$, we check if it matches the current target $t$ from $T$. If the current $t$ is reached, $t$ is updated to the next target in $T$. Subsequently, the $f$-cost of every state in the current $O$ is re-computed, where the $f$-cost in LLM-A* is computed as the sum of the state’s cost, the heuristic value, and the cost from the state to current $t$ (see in Algorithm 1:$20$), defined as $f(s)=g(s)+h(s)+cost(t,s)$. This step introduces an additional computational amount to the pathfinding process, and the time complexity scales linearly with both the length of $T$ and the increasing size of $O$. However, it is important that this re-computation process ensures that the $f$-cost of visited states in $O$ remains accurate and updated with the new target $t$.
@@ 4.1 Dataset
These parameters define the structure and constraints of each map, ensuring consistency and relevance to the standard experimental environment conditions for search-based algorithms. Notably, the discretization of points and actions within this continuous framework is a necessary simplification that allows us to make the problem tractable and effectively evaluate algorithms. Meanwhile, the map environment is able to scale properly for scalability experiment.
@@ 4.2 Experimental Setup
Our system facilitates search-based pathfinding within a scalable framework designed for environments of varying complexity. It consists of modules for environment management, agent control, and visualization (see Section 4.1).
  • Environment Management: Configures the environment and provides feedback, ensuring a challenging setup.
  • Agent Control: Customizes the agent’s operations using the algorithm and model, operating on discrete points and actions to make the problem tractable.
  • Visualization: Offers real-time and final visual outputs for comprehensive analysis.
While the environments considered are presented within a continuous framework, both the LLM and A* algorithms operate on discrete points and actions. This necessary simplification allows us to effectively evaluate the proposed LLM-A* algorithm, ensuring the system remains applicable to a wide range of complex environments.
Methodology Base Model Prompt Approach Operation Ratio $\downarrow$ (%) Storage Ratio $\downarrow$ (%) Relative Path Length $\downarrow$ (%) Valid Path Ratio $\uparrow$ (%) A* - - 100 100 100 100 Dynamic WA* $(w=2)$ 60.91 78.53 100.24 100 LLM GPT-3.5 Few-Shot - - 119.38 12.80 CoT - - 151.73 15.20 RePE - - 183.87 7.80 LLAMA3 Few-Shot - - 111.05 12.60 CoT - - 114.89 12.00 RePE - - 138.32 16.40 LLM-A* (Ours) GPT-3.5 Few-Shot 57.39 74.96 102.44 100 CoT 69.50 83.65 102.54 100 RePE 85.47 96.53 102.41 100 LLAMA3 Few-Shot 44.59 64.02 102.47 100 CoT 47.60 66.27 102.46 100 RePE 64.08 80.19 102.54 100
Table 1: Quantitative analysis of three pathfinding methodologies: the classical A* algorithm, dynamic weighted A* with initial weight of 2 and 0.99 decay, an LLM-only approach, and our proposed LLM-A* approach. The methodologies are evaluated on the map size ($50\times 30$) of original samples. The LLM-only approaches explore the path without explicitly searching the space grid by grid, so we do not report the operation and storage ratio. The table includes the results from GPT-3.5 and LLAMA3 models with three prompting approaches: Few-Shot, Chain of Thought (CoT), and Recursive Path Evaluation (RePE) for both LLM-only and LLM-A* approaches (see Section 4.4 for details).
@@ 4.4 Quantitative Analysis
When compared to other variants using non-admissible heuristics, LLM-A* demonstrates an superior performance in term of operation and storage efficiency. Dynamic weighted A* (with initial weight of 2 and 0.99 decay) employs a static logic to dynamically update the weight, but it lacks the flexibility inherent to LLM-A*. Consequently, dynamic weighted A* falls short in terms of both operation efficiency and storage efficiency compared to our advanced LLM-A* approach.
@@
The Recursive Path Evaluation (RePE) prompting method achieves marginal improvements in relative path length for the LLM-A* approach using GPT-3.5, with an increment of $2.41\%$. This suggests some potential of RePE’s step-by-step progression and intrinsic reasoning capabilities in generating more optimal waypoints, resulting in slightly more efficient paths. However, it is important to acknowledge that RePE underperforms compared to Chain of Thought (CoT) and few-shot prompting in a both operation and storage ratio, as well as efficiency in the LLM-only approach. This observation aligns with the limitations of LLMs in executing end-to-end path planning and spatial-temporal reasoning, which can affect their proficiency in sequentially reasoning out detailed path sequences and lead to issues such as hallucinations and misunderstandings, highlighting the diminish of RePE’s efficiency for long-horizon tasks, where the intermediate points chosen using this technique are not optimal. Therefore, while RePE shows some promise, its overall effectiveness is limited compared to other methods in both LLM-A* and LLM-only scenarios.
@@ 4.5 Qualitative Analysis
From the visualization in Figure 1, LLM-A* identifies the optimal path with only 140 operations, less than one-fifth the 859 operations required by A*, as well as the storage reduction. Both algorithms utilize a priority queue that stores the $f$-cost of each reached state, with the state having the lowest $f$-cost selected for exploration. The fundamental distinction between the two algorithms lies in their calculation of the $f$-cost or heuristic values. In addition, as the map size increases, this operational efficiency difference could become more pronounced, as further illustrated in Figure 3.
@@ 5 Conclusion
Although around 90% of the paths generated by LLM-A* are optimal, our algorithm does not guarantee optimal path. While these cases are relatively few, they indicate that the algorithm may sometimes yield paths that are not the shortest or most efficient. Future improvements could focus on enhancing the optimality of the generated paths to ensure more consistent performance.
Our experiments mainly utilized GPT-3.5-TURBO and LLAMA3-8B-16bit with basic prompt techniques. Although these models and prompts were adequate to validate the robustness of the LLM-A* algorithm, we did not explore a wider array of models or advanced prompt engineering strategies. Further testing with additional models and varied prompting methods could provide more comprehensive insights into the algorithm’s performance across different scenarios.
Acknowledgement
The work is partially support by a DARPA ANSR program FA8750-23-2-0004 and a National Science Foundation CAREER award #2339766. The views and conclusions are those of the authors and should not reflect the official policy or position of DARPA or the U.S. Government.